OpenStudy (anonymous):

solve for the given variable: t = 5s + 15 ; s

OpenStudy (anonymous):

t-15/s

OpenStudy (anonymous):

t-15/5

OpenStudy (anonymous):

i have to show work. i have no idea how you got that answer.

OpenStudy (anonymous):

well it is one equation and two unknowns hence the only way you can show the sol is 5s=t-15 s=(t-15)/5

OpenStudy (anonymous):

Oh, thank you for your help.

OpenStudy (anonymous):

would you do the same for w = uv + t ; u ?

OpenStudy (anonymous):

yes whatever you need, just isolate that variable